OpenStudy (purple_pink):

Madison bought a gift for $17.55. Sales tax of 8% was included in the price she paid. What was the price of the gift before the tax was added? A. $15.75 B. $16.15 C. $16.25 D. $16.47

OpenStudy (anonymous):

One second

OpenStudy (calculusxy):



OpenStudy (anonymous):

oh well @calculusxy looks like he is helping lol

OpenStudy (calculusxy):

100% represents the full price without the sales tax 108% representes the full price with the sales tax

OpenStudy (calculusxy):

@CupcakeQueen10 I am a she :)

OpenStudy (anonymous):

i didn't mean to put he sorry

OpenStudy (purple_pink):

oh lol

OpenStudy (calculusxy):

100% (full price) is what we don't know (\(x\)) 108% (full price + tax) is what we what do know because it says that the price with the sales tax of 8% is $17.55

OpenStudy (calculusxy):

We can use a proportion to solve this problem. \(\large \frac{100}{x} = \frac{108}{17.55}\)

OpenStudy (calculusxy):

Solve for \(x\).

OpenStudy (anonymous):

oh yea now i remember this is easy! lol

OpenStudy (purple_pink):

ok lemme calculate that

OpenStudy (anonymous):

I would do it but i am not supposed to

OpenStudy (anonymous):

lol i have to go sorry!! :D Bye!

OpenStudy (purple_pink):

ok the answer is x=16.25

OpenStudy (calculusxy):

Correct!

OpenStudy (calculusxy):

And if you want to check your work what you could do is: 1) 16.25 x 0.08 = ? ? + 16.25 should equal to 17.55 2) 16.25 x 1.08 which should equal to 17.55

OpenStudy (anonymous):

\[\frac{17.55}{1.08}=16.25 \]